How much acid must be added to 40ml of a 20% solution to increase it to a 28% solution?
Answer by chitra(359)

assuming 100% acid....

{100% }* x + [ 20% ]* (40 ml ) = ( 28 %) ( 40 + x )

where x = amount of acid in ml....

40 + x = total ml of 28% solution ...

convert % to decimal.... [ move decimal 2 times left on each ...]
1x + .20 * 40 = .28 ( 40 + x)

x + 8 = 11.2 + .28x

.72x = 3.2 ..... x = 4.44 ml